OpenStudy (anonymous):

I REALLY NEED HELP AND FAST. IF YOU KNOW MATH THAN PLZ HELP!!!! I WILL AWARD MEDALS!!!

OpenStudy (anonymous):

Given Trinagle ABC with A(-4, -2), B(4,4), and C(18,-8), answer questions 21-24. 1. Write the equation of the line containing the altitude that passes through B in standard form. 2. Write the equation of the line containing the median that passes through point C in slope-intercept form. 3. Write an equation for the line containing a perpendicular bisector of AC in point-slope form.

OpenStudy (anonymous):

Mostly need help on 2 and 3

OpenStudy (anonymous):

@SeaTurtle113 @FutureMathProfessor @Hero

OpenStudy (anonymous):

@Rea201 @lolo5656 @Hero @Kristen17 @Derrickhitch

OpenStudy (anonymous):

@hayhay5270

OpenStudy (anonymous):

I really need help with this 2

OpenStudy (anonymous):

you have the same question?

OpenStudy (anonymous):

yes the same 3

OpenStudy (anonymous):

well, does anyone have the answers? I know what number 1 is.

OpenStudy (anonymous):

could you please tell me one because I don't know that one

OpenStudy (anonymous):

What's the "median"?

OpenStudy (anonymous):

please but explain 2

OpenStudy (anonymous):

Cab you tell me number 2?

OpenStudy (anonymous):

no I don't understand any of this

OpenStudy (anonymous):

well number 1 I used Desmos graphing calculator and I figured it out a little bit, so I went from there.

OpenStudy (anonymous):

but like what did you get

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

do you even know how to do this?

OpenStudy (anonymous):

no I dont

OpenStudy (anonymous):

Just find the slope and put it into y=mx+b. All you need is the slope(m) and the y-intercept, whcih is given.

OpenStudy (anonymous):

the slope is 27/100 that is what I got

OpenStudy (anonymous):

Well, if no one knows the answer or know how to get to them, I am going to close the question.

OpenStudy (anonymous):

no please help me with one

OpenStudy (anonymous):

I don't think that is the right slope but I keep getting that

OpenStudy (anonymous):

@Smarts please help please please please

OpenStudy (anonymous):

Ill give you this, the slope is 3.4. At least for line AB it is.

OpenStudy (anonymous):

so it would be y=4/3x=?

OpenStudy (anonymous):

I meant 3/4

OpenStudy (anonymous):

sorry y=3/4x+what

OpenStudy (anonymous):

Thats the ewasy part:) Ur welcome:)

OpenStudy (anonymous):

I mean 4/3 cause its perpendicular right

OpenStudy (anonymous):

the slope of the line is 4/3 right

OpenStudy (anonymous):

@Smarts

OpenStudy (anonymous):

please smarts help please just with the one please I really need help

OpenStudy (anonymous):

@smarts please help I would help you if I could your a hipacrite if you don't help because you wanted people to help you and nobody is if you cant eve help somone

OpenStudy (anonymous):

@Smarts